Diffusion Morning Report 03/29/04

ted berg tberg at snf.stanford.edu
Mon Mar 29 08:30:59 PST 2004


Good Morning all,
    Looking pretty good this morning. Had a temp issue with Nitride, 
remade tc connections on profile and tightened spikes.  Temp is falling now.
SiGe is getting closer  on Temp Status. Everything else looking Ok

Diffusion 	Date: 	3/29/2004 	 
Tube Name/Number 	Status 	Comments 	THICKNESS-Change at
Tylan 1 	Up 	  	 
Tylan 2 	Up 	  	 
Tylan 3 	Up 	  	 
Tylan 4 	Up 	  	 
Tylan 5-BBr3 	Up 	  	 
Tylan 6-POCl3 	Up 	  	 
Tylan 7-TEOS2 	down 	Waiting for Mfc-3/26 delivery 	 
Tylan 8-SiGe 	down 	Temp cal 	5.4/30
Tylan 9-Poly 	Up 	  	13.1/30
Tylan 10-Nitride 	yellow 	particles-temp issues 	6.25/30
Tylan 11-BPSG-LTO 	Up 	  	8.9/30
Tylan 12-FGA 	Up 	  	 
Tystar 	Up 	  	1.2/30
Doped Poly 	GONE 	  	 
TEOS 	yellow 	 running 	 
Tylan 16 	N/A 	  	 
Burn Box 	Up 	  	 
ZYGO 	Up 	keyboard replaced 	 
  	  	  	 
GAAS 22 	UP 	  	 
GAAS 23 	UP 	  	 
  	  	  	 
SiH4 for furnaces 	  	Auto changeover working well 	A on line 640 PSI


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20040329/9d27e5ed/attachment.html>


More information about the diffusion mailing list